OpenStudy (dnel1):

What is the empirical formula of a compound that contains 43.38% sodium, 11.33% carbon, and 45.29% oxygen?

OpenStudy (cuanchi):

Na2CO3

OpenStudy (kevin):

You need to find its mole ratio assume 1% = 1 gram Then it will be :  \[\frac{ 43.38 }{ 23 } : \frac{ 11.33 }{ 12 } :\frac{ 45.29% }{ 16 }\] You will got the smallest mole ration : 2 : 1 : 3 So, the empiric formula will become Na2CO3